OpenStudy (anonymous):

Add 1 10 + 11 100 and select the correct answer. A) 1 100 B) 1 5 C) 11 100 D) 21 100

OpenStudy (anonymous):

these are fractions

OpenStudy (anonymous):

\[\frac{ 1 }{ 10}+\frac{ 11 }{ 100 }\]

OpenStudy (anonymous):

like this

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first you look at denominator so if u do 10 times 10 u get 100 right

OpenStudy (anonymous):

so u need to make the denominator of both equal to 100 and to do the term above 10 must be multiplied by 10

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

\[\frac{ 1 \times 10 +11 }{ 100 }\]

OpenStudy (anonymous):

\[\frac{ 21 }{ 100 }\]

OpenStudy (anonymous):

ok thank you so much You are a really good helper

OpenStudy (anonymous):

so the answer is D right

oregonduck (oregonduck):

yes d

OpenStudy (anonymous):

Yep! The answer is D

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

fanned